4/9*n=1 1/3
I know that I divide and carry over 4/9. Do I flip the fraction as I take it over to the 1 1/3 side?

Answer by Cromlix(4381)

4/9 * n = 1 1/3
1 1/3 = 4/3
4/9 * n = 4/3
9*4/4*3 = n
36/12 = n
n = 3
Hope this helps.
:-)